OpenStudy (anonymous):

Find the limit and use L'Hospital's Rule when appropriate. limit of (5^t - 3^t)/t as t approaches 0

OpenStudy (dumbcow):

differentiate top and bottom: \[\rightarrow \lim_{t \rightarrow 0} \frac{\ln(5) 5^{t} - \ln(3) 3^{t}}{1} = \ln(5) -\ln(3)\]

OpenStudy (anonymous):

agreed

OpenStudy (anonymous):

Thanks! I forgot that the derivative of a^x = a^x * In a :)

OpenStudy (dumbcow):

:)